OpenStudy (anonymous):

A plane flies round-trip to Philadelphia. It flies to Philadelphia at 220 miles per hour and back home with a tailwind at 280 miles per hour. If the total trip takes 6.5 hours, how many miles does the plane fly round-trip?

OpenStudy (anonymous):

we want distance, so lets make \(D\) = distance now lets see if we can get an expression involving time, because that is the number you know \(T=6.5\)

OpenStudy (anonymous):

time is distance divided by rate i.e. \(T=\frac{D}{R}\) the time going is \(\frac{D}{220}\) and the time returning is \(\frac{D}{280}\) so the total time is \[\frac{D}{220}+\frac{D}{280}\] which you know is \(6.5\) solve \[\frac{D}{220}+\frac{D}{280}=6.5\] for \(D\)

OpenStudy (anonymous):

i still dont understand

OpenStudy (anonymous):

that is my best explanation i do not have another one, sorry

OpenStudy (anonymous):

its okay thanks though